OpenStudy (anonymous):

Solve the SYSTEM please help!! x = y + 11 2x + y += -2 A. no solution b. (-13, 2) c. (0, -11) d. (3, -8)

OpenStudy (igreen):

There are many different ways to solve systems. For this one, the best one is substitution. Substitute 'y + 11'(from the first equation) for x in the second equation.

OpenStudy (anonymous):

i'm still confused..

OpenStudy (igreen):

2x + y = -2 Input 'y + 11' for x: 2(y + 11) + y = -2 Distribute 2 into the parenthesis: 2y + 22 + y = -2 Subtract 22 to both sides: 2y + y = -24 Add: 3y = -24 Can you solve the rest? @kaileyperez11

OpenStudy (anonymous):

so, the answer is (3, -8) correct? @iGreen

OpenStudy (igreen):

Yes!

OpenStudy (anonymous):

ok thanks so much!

OpenStudy (igreen):

No problem!